### v2

# The StudyChat Dataset: Student Dialogues With ChatGPT in an Artificial Intelligence Course

Abstract
The widespread availability of large language models (LLMs), such as ChatGPT, has significantly impacted education, raising both opportunities and challenges. Students can frequently interact with LLM-powered, interactive learning tools, but their usage patterns need to be analyzed to ensure ethical usage of these tools. To better understand how students interact with LLMs in an academic setting, we introduce StudyChat, a publicly available dataset capturing real-world student interactions with an LLM-powered tutoring chatbot in a semester-long, university-level artificial intelligence (AI) course. We deploy a web application that replicates ChatGPT’s core functionalities, and use it to log student interactions with the LLM while working on programming assignments. We collect 1,197 conversations, which we annotate using a dialogue act labeling schema inspired by observed interaction patterns and prior research. Additionally, we analyze these interactions, highlight behavioral trends, and analyze how specific usage patterns relate to course outcomes. StudyChat provides a rich resource for the learning sciences and AI in education communities, enabling further research into the evolving role of LLMs in education  [^1] https://huggingface.co/datasets/wmcnicho/StudyChat.
Keywords Conversational Datasets  $\cdot$ CS Education  $\cdot$ Dialogue Analysis  $\cdot$ Large Language Models  $\cdot$ Natural Language Processing
1 Introduction
Advances in large language models (LLMs) have had major ramifications on almost all levels of education. Many LLM-powered educational technologies have recently been developed, leveraging the generative capabilities of these models to create new course content (MacNeil et al., 2022), assignments (MacNeil et al., 2023), and interactive learning experiences(Jin et al., 2024; Taneja et al., 2024). These interactive technologies enable rich engagement with course material by allowing students to engage in on-demand conversations with an interactive tutor, and receive personalized feedback on their work. However, while the potential use cases for LLMs in education are vast, many educators are understandably concerned about student misuse of LLM chatbot tools, such as ChatGPTLau and Guo (2023). While a motivated learner can use these tools in a learning positive way, an unmotivated student could use them to complete their assignments on their behalf, circumventing the learning objectives of the assignment, or become over-reliant on the model(Bastani et al., 2024).
This dichotomy between learning opportunity and academic risk necessitates an exploration into understanding student behavior with LLM chatbot tools. A more nuanced understanding of this behavior could inform further development of LLM-based Intelligent Tutoring Systems (ITS) that dynamically adapt to students, encouraging learning-focused behaviors and discouraging counterproductive ones. Furthermore, a deeper understanding of student behavior can lead to more robust student models, which better estimate knowledge states, misconceptions, and learning trajectories. For example, such a model could potentially detect when a student is overly reliant on LLM responses and flag this to instructors in real-time, or be used to help develop assessments which are specifically designed to test for this over-reliance.
In this work, we conduct a case study to explore student behavior with an LLM, specifically in the context of upper-division, university-level Computer Science (CS) education. We develop a web application with the same core features and underlying LLM as ChatGPT, and deploy it in an introductory-level course on artificial intelligence (AI). We encourage students to use this tool, without restrictions, on programming assignments over the course of the semester. With student consent, we recorded over 1,000 student-LLM conversations from 84 students across 7 unique assignments, completed over 15 weeks. We also collect 455 assignment submissions from 65 students who further agreed to share their work. In addition, we annotate these conversation turns with a dialogue act labeling schema; the schema is inspired by other existing studies on co-programming behavior and our observations during a systematic review of the conversations. We leverage an LLM prompting-based approach to perform this annotation at scale and validate this approach with a human evaluation.
To explore how students engage with an LLM during programming assignments, we analyze these dialogue act labels and highlight notable trends in behavior. In addition, we investigate if certain patterns of behavior correlate with course performance by performing a regression analysis. Our key findings are 1.) students primarily use the LLM to ask questions, such as assignment clarification or code explanation, but co-programming and answer seeking behaviors are also very common 2.) certain engagement patterns such as asking conceptual questions or actively discussing problem-solving approaches with the LLM, correlate with better scores on graded items, and 3.) students who tend to use the LLM for debugging and tool assistance tend to perform worse on tool-prohibited items which perhaps highlights LLM over-reliance or under-preparedness. These findings provide data-driven insights into how students interact with LLMs. We release this to the broader research community to foster collaboration and exploration into LLM’s impact on CS education.
2 Related Works
2.1 ChatGPT Classroom Studies
Since its release in November 2022, there has been a vast amount of research interest into how ChatGPT can be utilized in classroom settings. Studies have explored various generative applications of ChatGPT for course exercises (Xiao et al., 2023), code explanations (Chen et al., 2023), multiple-choice distractors (Feng et al., 2024), and more (Khan et al., 2023; Li et al., 2024; Hsieh and Yang, 2024). Direct classroom studies have also explored how ChatGPT-powered applications can answer student questions as a virtual teaching assistant (Taneja et al., 2024), help students learn by teaching chatbots to code (Jin et al., 2024), and even provide coaching and feedback to instructors on content delivery(Wang and Demszky, 2023).
Closely related to this work are a handful of cases studies exploring student use of ChatGPT in university level CS classrooms. Ghimire et. al explore student use of ChatGPT for two assignments in a CS1 programming course, finding students often ask questions of the system instead of directly seeking solutions (Ghimire and Edwards, 2024). Ma et. al also perform a case study where CS1 students were allowed to use ChatGPT freely on in-class exercises and on a course project, but only for conceptual questions on homework(Ma et al., 2024). Additionally, Brender et. al perform a case study on how graduate students in a robotics course utilize ChatGPT on a voluntary assignment, controlling for learning outcomes and clustering behavior into broad categories (Brender et al., 2024). Analysis of student prompts in these studies have noted a student trend to use models to ask questions more than directly writing code solutions which is consistent with our observations and those by Bhalerao (2024).Surveys on student perceptions towards LLMs have also been conducted which suggest that CS student adoption of LLMs is nuanced and not universal(Weber et al., 2024). Similarly, we observe that a some students use the system heavily while most students engage moderately with the model, revealing a range of adoption behaviors.
2.2 Education Discourse Analysis
It has been demonstrated by many that tutoring is a highly effective form of instruction Cohen et al. (1982). Thus, human tutoring dialogue has been studied extensively Graesser et al. (1995). In recent years, dialogue act (DA) classification on conversation turns is often used in this discourse analysisChen et al. (2011). DA labels classify a participant’s utterance by the function it serves in the conversation. When DA labeling is applied to an entire corpus of dialogues, it enables discourse-level analysis into patterns of conversation that are most frequent or significant. In an education context, significant patterns are often those associated with better learning outcomes, as such existing works have studied such patterns in tutor-student or student-student learning environmentsGraesser et al. (1995); Earle-Randell et al. (2023). Existing works also analyze human-machine classroom conversations between students and ITS chatbots Vail and Boyer (2014); Graesser et al. (2004); Fisher et al. (2020). These works provide promising insights which can inform more effective tutoring pedagogyRus et al. (2017) and interactive learning toolsChen et al. (2011).
While DA analysis is an attractive approach to discourse analysis, the labeling process poses many challenges, solutions to which constitute an large body of workBoyer et al. (2011); Tan et al. (2023); Ezen-Can and Boyer (2015); Lin et al. (2023); Boyer et al. (2010); Kumaran et al. (2023). First, there exists the task of annotating a large corpus of conversations. To automate such a task, various machine learning based approaches to this have been used such as Hidden Markov ModelsBoyer et al. (2010), decision treesSamei et al. (2014), and Markov Random FieldsEzen-Can and Boyer (2015). In our work, we employ a LLM prompt-based approach to dialogue act labeling, similar to that ofWang et al. (2023). Second, there exists the challenge of robust schema design, where existing work has looked into refining prior labeling schemata to be less ambiguousVail and Boyer (2014) and to try to develop classroom agnostic labeling schemaHennessy et al. (2016).
2.3 Education Dialog Datasets
In an effort to collectively improve data-driven learning technologies, there have been recent efforts to expand the amount of education dialog datasets in a variety of learning domains. Datasets such as TalkMoves and NCTE directly capture and annotate classroom discourse in K-12 lessons Suresh et al. (2022); Demszky and Hill (2022). Other works employ crowd workers or leverage LLMs to simulate the student or teacher role, such as MathDialMacina et al. (2023), Teacher-Student Chatroom Corpus (TSCC)Caines et al. (2020), Conversational Instruction with Multi-responses and Actions (CIMA)Stasaski et al. (2020), and GrounDialogZhang et al. (2023). Our work contributes to this growing body of research by providing a large-scale, annotated dialog dataset of in a university-level CS course, offering insights into how students engage with LLM-chatbot tools in a real learning environment.
Figure 1: The web application used in this study. The user interface mirrors functionality of ChatGPT and allows students to organize conversations by assignment.
3 Dataset
In this section, we detail the process by which the StudyChat dataset was formed via the classroom context, application, and dialogue act annotation process.
Table 1: Dialogue act schema for student interactions with ChatGPT.
Category Acts Description
Writing Write Code, Write English, Conversion, Summarize, Other The user asks the LLM to generate text, such as code, explanations, or summaries.
Editing Edit Code, Edit English, Other The user asks for modifications or improvements to existing code or text.
Contextual Questions Assignment Clarification, Code Explanation, Interpret Output, Other The user asks a question based on something from the assignment or prior conversation.
Conceptual Questions Programming Language, Python Library, Computer Science, Programming Tools, Other Concept The user asks a general question about programming, libraries, tools, or CS concepts unrelated to the conversation.
Verification Verify Code, Verify Report, Verify Output, Other The user requests validation of their code, writing, or output to ensure correctness.
Context Assignment Information, Error Message, Code, Other The user supplies relevant information to the LLM, such as assignment instructions, code snippets, or error messages.
Off Topic Chit-Chat, Greeting, Gratitude, Other Conversations unrelated to the assignment, such as casual discussions, greetings, or expressions of gratitude.
Misc Other A catch-all category for dialogue acts that do not fit elsewhere.
3.1 Participants and Course Context
Our IRB-approved study was conducted among students aged 18 or above, who were enrolled in an AI course at a US R1 university in the Fall 2024 semester. 117 students were enrolled in the course, of which 84 agreed to participate in the study. Students who did not participate were also given equivalent access to use the tool over the semester, but their data was not collected. Each week, students attend two 75-minute lectures and independently complete assignments related to the topics covered in lecture. The course consists of 3 modules taught over 15 weeks. During each module, students complete 2-3 Python programming assignments related to the lecture material. In addition to the assignments, students were assessed via 2 end-of-module exams, a final exam, and a group project.
3.2 Data Collection
In order to effectively collect student interactions, we developed and deployed a web application designed to function similar to consumer LLM chatbot interfaces, as shown in Figure 1. As our backing chatbot LLM, we used gpt-4o-mini, the LLM for the free tier of ChatGPT at the time of this studygpt . Students were instructed they could use this ChatGPT-based tool, without restrictions, on all assignments in the course and the final project. It was emphasized that their approach to using the tool would not affect their grade and they were encouraged to use in any manner of their choosing. We used the default "You are a helpful assistant" system prompt, which does not discouraging answer-seeking behavior, and did not provide the chatbot with any course materials for reference. The reasoning for these decisions was to encourage students to use our platform instead of seeking an outside application for homework help and that the collected conversations would more accurately reflect how students generally use non-instructional, consumer chatbots. In addition to the conversations, we also collected the 455 submissions by 65 of the 84 student participants consented to share their assignment solutions. Finally we collected responses to a pre-study survey on students’ perspectives on LLM-based programming tools.
3.3 Filtering of Student Identifiable Information
To remove any Personally Identifiable Information (PII) from the conversations, we systematically analyze the conversations to uncover patterns of PII exposure. We first randomly sample 12 conversations from each course topic and read through the conversations, scanning for traces of PII. Our main finding of potentially sensitive student information is in the form of pasted terminal outputs and error messages, which sometimes contain PII, such as a user’s directory or GitHub account name. Compiling these common patterns, we then develop a regular expression-based script to remove these patterns and any direct matches to known user PII in the course roster. Across the entire dataset, this process removes 2009 cases of possible PII exposure. Then, we sample 12 additional conversations from each topic and performed the same procedure. In our second pass of the second set of conversations, we did not find any additional patterns or cases of PII exposure. We acknowledge that although seemingly effective, this process does not guarantee the removal of all PII exposures.
Figure 2: Violin plot showing the distribution of total utterances per user for each assignment topic. The width of each violin represents the density of user interactions, while inner quartiles indicate the spread of engagement.
3.4 Data Statistics
Across all conversations, we collect 7,856 student utterances and corresponding LLM responses over 7 graded assignments. The mean conversation length in the dataset is 7.07 utterances, with a median of 3 utterances per conversation session. On average, each student has 23.13 utterances and 13.5 distinct conversations with the chatbot. Figure 2 shows the number of utterances for each user across topics, where there is an average of 70.89 utterances for each topic. We see a skewed distribution where a small group of students use the model extensively, over 100 utterances for some assignments, while the average median number of utterances per student is 7.7. Specifically, we find that the top 5% most-active students account for 21.82% of all utterances in the dataset.
3.5 Dialogue Act Annotation
In order to better understand common behavior and discourse-level trends in the conversations, we develop a dialogue act labeling schema for student turns of the conversations. The schema, as shown in Table 1 is two-leveled and captures common user behaviors in the dataset. This scheme is developed on top of existing work Wood et al. (2018) which conducted a “Wizard of Oz” experiment with programmers who were debugging alongside a simulated virtual assistant. It also mirrors our own observations after systematically reviewing the conversations, and aligns with those of Ma. et alMa et al. (2024).
First, we sample a set of 150 utterances and develop an initial schema with labeling guidelines, including examples and descriptions for each dialogue act. We then provide them to a research assistant, who independently labels the same set given our initial schema. We iterate on our labeling instructions and schema by computing the agreement between the two annotators and identifying sources of disagreement. After this iteration, our inter-rater agreement reached a Cohen’s kappaCohen (1960) of 0.910 for broad level labels and 0.788 for specific labels, indicating almost perfect agreement on the broad category of dialogue act and substantial agreement on the specific label Landis and Koch (1977). To validate this labeling, we provided these final instructions to two additional annotators who independently label the same set of utterances. We observed and average 4-way inter-rater agreement Cohen’s kappa of 0.740 for broad level agreement and 0.575 for specific level agreement, indicating substantial agreement on broad level labeling and moderate agreement on specific labeling.
To apply this labeling at scale to the reminder of the dataset and since manual labeling is extraneous, we perform an LLM-prompting based approach similar to prior workWang et al. (2023). We prompt the LLM with detailed instructions on the task, the annotation instructions, and the entire conversation up until the student turn to be labeled, and ask the LLM to provide the label based on our schema, along with a justification. We utilize GPT4o to perform this labeling, which we applied to the entire PII-filtered dataset.
To validate the LLM labels, we again systematically sample 154 utterances, distinct from those used in developing the initial labeling schema. We provide them to the same 4 human annotators without any further discussion or instructions. The resulting human inter-annotator agreement on the held out set is 0.578 on broad labels and 0.538 on specific labels, indicating moderate agreement between human annotators on the held out set. This drop in agreement is expected, since the new utterances in the held out set present new ambiguous utterances to which multiple labels may apply. The average human-LLM agreement on the held out set is 0.446 on the broad labels and 0.333 on the specific labels, indicating moderate and fair agreement respectively. There are many potential causes for the disparity between human and LLM labeling; the most likely reason is that the LLM does not have knowledge of the course context like the human annotators, who have either taught the course or completed it as a student in prior semesters.
Overall, we emphasize that there are challenges in creating a universal dialogue act labeling schema for any domain, and acknowledge that there is room for further refinement, which we leave for future work since it is not the focus of the current paper, and is the subject of many related works Boyer et al. (2011); Tan et al. (2023); Ezen-Can and Boyer (2015); Lin et al. (2023); Boyer et al. (2010); Kumaran et al. (2023).
4 Results and Analysis
In this section, we report on the overall trends we observe in the StudyChat dataset. We also analyze how students’ interactions with ChatGPT relate to course performance via a regression analysis.
Table 2: Distribution of dialogue act labels in StudyChat.
Conceptual Questions Contextual Questions Writing
Programming Language 219 Assignment Clarification 561 Write Code 617
Python Library 373 Code Explanation 783 Write English 163
Computer Science 113 Interpret Output 74 Conversion 207
Programming Tools 98 Other 1882 Summarize 64
Other Concept 352 Other 380
Total 1155 Total 3300 Total 1431
Verification Context Editing
Verify Code 237 Assignment Information 197 Edit Code 183
Verify Report 49 Error Message 253 Edit English 103
Verify Output 84 Code 268 Other 0
Other 44 Other 311
Total 414 Total 1029 Total 286
4.1 Dialogue Act Analysis
Table 2 shows the distribution of dialog acts in StudyChat. Off-Topic (173 counts) and Misc (68 counts) categories are relatively infrequent and not included in the table. The most common categories of dialogue acts are Contextual Questions, specifically the Assignment Clarification and Code Explanation acts (excluding than the generic dialogue act of Other). The relatively high frequency of contextual questions suggests that students are usually using the LLM to engage with and better understand the assignments, instead of simply copy-pasting LLM solutions. These observations are somewhat similar to those in Ma et al. (2024): students mostly tend to ask questions to the LLM rather than using it to directly complete assignments on their behalf. However, in our study, students were not discouraged from using the LLM to help co-author their code which may explain why in our study Writing > Write Code is still a very frequently occurring dialogue act.
4.2 Course Outcome Regression Analysis
We now investigate the relationship between student interactions with ChatGPT and their course performance via a multiple linear regression analysis. Our primary goal is to compare how different types of student-LLM interactions correlate with scores on items where ChatGPT was allowed and items where it was prohibited.
4.2.1 Feature selection
We select independent variables from two sources, the dialogue act labels and a pre-assessment survey given to students at the beginning of the course. We normalize all 42 total features against the maximum value across students, resulting in feature values in the range $[0,1]$. The dialogue act features are simply the counts of each type of dialogue act across each student’s utterances, and the student’s total number of utterances. In the survey, we ask students a series of questions about their background and prior experience, such as whether or not they had an prior experience with AI or machine learning. Then, on a 5-point Likert scale, we ask students to report: their confidence in three pre-requisite areas, including Python, Discrete Math, and Probability; how often they use LLM tools (from “Never” to “Everyday”); and how important they think LLM tools help their success in computer science.
Table 3: Regression coefficients for features with p-value $<0.2$.
Category Feature $\mathbf{\beta}$ p-value
ChatGPT Allowed
Provide Context Assignment Information 0.2249 0.03903
Conceptual Questions Other Concept 0.1111 0.04452
Contextual Questions Python Library 0.1178 0.15625
Verification Other -0.1366 0.18216
ChatGPT Prohibited
Provide Context Error Message -0.2937 0.02884
Provide Context Other 0.1904 0.05377
Pre-Survey Question Discrete Math Confidence 0.1557 0.06331
Conceptual Questions Programming Tools -0.1441 0.06613
Writing Request Code/Data Conversion 0.1247 0.06963
Contextual Questions Programming Language 0.1377 0.08329
Editing Request Edit Code 0.3072 0.18720
4.2.2 Regression Results
We fit two separate multiple linear regression models: The first model predicts student performance on the assignments and final group project, where using ChatGPT is allowed. The second model predicts student performance on quizzes and exams, where ChatGPT is prohibited. The dependent in both models is the aggregated score, normalized across all students in the dataset, also in a range $[0,1]$. We acknowledge that the StudyChat dataset only includes the scores of a subset of students who consented to share their assignment submissions, which somewhat limits the reproducibly of these results.
We observe an $R^{2}$ of 0.4136 on scores where ChatGPT usage is allowed and an $R^{2}$ of 0.5031 on scores where it is prohibited. We detail the most statistically significant features for each regression in Table 3. For items where ChatGPT usage is allowed, two features are statistically significant (p-value $<0.05$): Context > Assignment Information and Conceptual Question > Other. Both are positively correlated with student performance. We review 50 samples of utterances with these labels to explore trends in student behavior. For the Context > Assignment Information label, we find that most of these samples are instances of the student directly inserting all or parts of the instructions from the programming assignment into the chatbot. Therefore, it is unsurprising that this behavior would correlate with higher scores on the corresponding assignment, since it indicates the students are directly seeing, and perhaps copying ChatGPT’s solutions to the assignment. On the other hand, when reviewing the Conceptual Question > Other samples, we observe interactions where students ask questions that are tangential to the course, on topics such as statistics, ethical computing, computer networking, C++ programming libraries, and more. One explanation for the correlation between this feature and student performance is that it is a proxy for student interest and general curiosity about computer science, which may translate to better course engagement and completion.
In contrast, for quizzes and exams where ChatGPT usage is prohibited, only one feature, Context > Error Message, is statistically significant and it negatively correlates with performance. This observation suggests that students who tend to use ChatGPT to debug and fix error messages perform worse on tests. This finding seems counterintuitive: the course examinations do not ask any Python or library related questions and are focused on AI concepts and algorithms, so a student’s use of an LLM to help debug compiler errors seems unrelated. One possible explanation is this observation indicates student over-reliance on ChatGPT in completing the programming assignments, which may result in a lack of engagement with the core algorithms and concepts present in the assignment.
Extending our observations to borderline significant features (p-value $<0.07$), we see that Conceptual Questions > Programming Tools is also negatively correlated with exam performance, which could further support the hypothesis above of student over-reliance on ChatGPT. On the other hand, there are a number of borderline features that correlate with positive performance, such as Context > Other. When investigating these utterances, we see that these utterances show active student engagement in co-programming with the LLM. In these cases, unlike in the Context > Assignment Information feature, which is direct copy-pasting, the student is actively describing the requirements of the assignment in their own words to provide further information to the LLM. This behavior indicates the student’s active understanding of the assignment and underlying concepts, as they articulate what is needed to make progress on the assignment with a conceptual understanding of what they are doing. To put these features in perspective, students’ self-reported confidence in discrete math has a clear connection to exam performance, since data structure and algorithm knowledge are important background to the concepts covered in exams. However, this confidence feature is just as explanatory as the dialogue features and similarly significant. Finally, the Writing > Conversion feature is difficult to explain. In one of the assignments, students are directly asked to convert a table from one data format to another with ChatGPT, in order to explore the text generation capabilities of LLMs. In our review of the data, we only see cases of students completing this part of the assignment and no other conversions, which only indicates that they are following the assignment instructions.
Our initial regression analysis of StudyChat provides some insights, albeit noisy ones, into the relationship between student behavior with LLMs and course performance. We see that students copy-pasting behaviors correlate with better assignment grades, but we do not observe a negative correlation between this behavior and exam performance, but instead see this correlation with help-seeking behaviors (tool assistance and debugging). These observations suggest that there could be student behavior patterns with LLMs that could potentially be monitored and detected in real time, enabling teachers to intervene and help students who are exhibiting patterns that correlate with at-risk performance.
5 Future Research Directions
In this work, we provide an initial analysis of the StudyChat dataset and surface-level insights into this data. In this section, we outline additional avenues for possible future work on this data and the kind of research questions that can be asked and possibly answered.
First, since this dataset includes time stamps on each student interaction with the LLM, we can track student behavior over time, both within an assignment and across assignments over the course of the semester. For example, we can study the variation in student engagement levels over each assignment: whether student activity drops off over the semester, and if there are any correlations between disengagement and course performance. In addition, we can study whether increased engagement with an LLM leads to better course performance and assignment solutions.
Second, since this dataset also includes student assignment submissions, we can further compare these submissions against student utterances when talking to the LLM. For example, we can study how closely final student submissions mirror the exact LLM output, to quantify whether students are simply copy-pasting or using LLM output for helpful suggestions. We can also investigate how such behavior correlates with course performance. This analysis may inform future content developers to develop homework questions that cannot be answered by solely relying on LLM output or perhaps integrate prompt strategies into the assignments that teach effective usage of LLMs and common pitfalls.
Third, since our initial labeling schema has clear room for improvement, we can investigate more accurate and fine-grained methods of classifying student utterances. For example, we can use span-level tagging to isolate parts of student utterance that are of particular interest, which may enable a more thorough discourse analysis and findings. Moreover, one may investigate n-gram dialogue acts and even study how student intent evolves over a conversation, which may reveal more insights into how students react to different LLM output. These insights can potentially inform future work on aligning LLMs with pedagogical principles.
Finally, our initial findings of dialogue acts which correlate with negative student performance is only a beginning into a proper investigation into the potential negative learning side-effects of LLM usage. LLMs are not oracles, and when they hallucinate information they could confuse students or worse, potentially instill misconceptions. A more thorough investigation into the accuracy of the statements made by the chatbot is certainly warranted, and if false statements made by the LLM lead to more buggy code or worse course outcomes.
6 Conclusions
In this work we present the StudyChat dataset, a collection of over 1,000 conversations between students and ChatGPT in an upper-division, university-level artificial intelligence course. We systematically create a dialogue act labeling schema which we applied to all utterances in the dataset via LLM labeling and validate this labeling with human agreement. We analyze the resulting labeled dataset via a regression analysis of student behavior, which suggests that students tend to use the model largely for question answering more than any other activity and, in some cases, this correlates to better course performance, but other help-seeking behaviors correlate with lower student performance.
Acknowledgments
We would like to thank all the students of this class for their participation and engagement in this project. We are also grateful to the entire course staff team for their hard work and dedication to teaching.
We extend a special thanks to Aditya Singh for his insightful discussions which helped shape this work, and for his contributions to the design and implementation of many course assignments.